# What Ails Generative Structure-based Drug Design: Too Little or Too Much Expressivity?

### v1

@@ 4.3 Atom sampler
To define the generative model over the atom types given unlabelled structure, we adapt an existing generative model MoFlow (Zang and Wang, 2020), consisting of an unlabelled graph model $f_{U}$ and a conditional atom model $f_{\mathbf{a}|U}$. The atom model induces a conditional distribution $p_{f_{\mathbf{a}|U}}(\mathbf{a}^{M}|U^{M})$, which we use as our atom sampling distribution:
@@
Table 1: SimpleSBDD finds high-quality drug candidates up to 1000x faster than competing methods. Comparison of multiple properties of the CrossDocked2020 test data and other generation methods. $\dagger$ denotes method run exclusively on CPU. "-" denotes that a metric was not reported and we were unable to estimate. Please refer to Appendix H for more details regarding the comparison.
@@ 5 Results
We follow related work of Peng et al. (2022); Schneuing et al. (2022) and use the CrossDocked2020 dataset (Francoeur et al., 2020) for evaluating the models. We use the same train-test split, which separates protein pockets based on their sequence similarity computed with MMseqs2 (Steinegger and Söding, 2017) and contains 100,000 train and 100 test protein-ligand pairs.
@@ 5.1 Assessing general properties of sampled molecules
We follow the evaluation procedure described by Peng et al. (2022). For each of 100 test protein pockets, we sample 100 molecules. As evaluation metrics we report 1) Vina Score estimating binding affinity (Alhossary et al., 2015), 2) High Affinity, which is the percentage of generated molecules with binding affinity at least as good as ground truth, 3) QED, a quantitative estimate of druglikeness, 4) SA, synthetic accesibility, 5) Diversity defined by average pairwise Tanimoto dissimilarity for the generated molecules in the pocket, 6) Novelty, defined by average Tanimoto dissimilarity to the most similar molecule from training set, 7) #Params, the number of trainable parameters and 8) Time to generate 100 molecules in seconds. We include more molecular metrics in Appendix G. As baselines, we use eight recent methods for SBDD: Pocket2Mol (Peng et al., 2022), DiffSBDD (Schneuing et al., 2022), TargetDiff (Guan et al., 2023a), DecompDiff (Guan et al., 2023b), FLAG (Zhang et al., 2022), DrugGPS (Zhang and Liu, 2023), D3FG (Lin et al., 2024) and EQGAT-diff (Le et al., 2024).
@@ 5.2 Additional evaluation
We have demonstrated that SimpleSBDD is very competitive on the standard evaluation criteria for SBDD models. We now discuss broader applications of our proposed model.
@@
Due to our novel decomposition of the model into the unlabelled graph and atom labels, we can optimize the estimated binding affinity and other properties simultaneously. This is especially important, since molecules without desirable drug-like properties are of significantly diminished practical utility even if the predicted binding is strong. Specifically, given the target protein we first generate an unlabelled graph with strong predicted binding affinity. As we noted in Section 4.2 and elaborated on in Appendix C, there is large diversity in the chemical properties of molecules that share the unlabelled graph. Therefore, for the sampled unlabelled graph, there is enough flexibility in atom types to enable molecular property optimization.
As an illustrative example, we follow Gómez-Bombarelli et al. (2018) and choose $5\cdot\mbox{QED}+\mbox{SA}$ as the property mixture to optimize alongside predicted binding affinity. Concretely, 50 proposals are sampled using the atom type model defined in Equation (7) and the candidate with the best property values is selected. We denote this model variant SimpleSBDD–$\mathcal{PO}$. With this additional procedure we can significantly improve all molecular metrics without sacrificing binding affinity whilst remaining an order of magnitude faster than baselines. See Appendix I for more details.
@@
So far we have compared our method with multiple baselines which generate molecules conditioned on a target protein. Neither SimpleSBDD nor any of these baselines has access to the Vina software during generation. We now also compare with optimization-based methods, i.e. algorithms, which generate molecules by explicitly using Vina during sampling to directly optimize it. Specifically, we compare SimpleSBDD–$\mathcal{PO}$ with the three best performing methods, i.e. Reinforced Genetic Algorithm (Fu et al., 2022), 3D-Monte Carlo Tree Search (3D-MCTS) (Du et al., 2023) and AutoGrow4.0 (Spiegel and Durrant, 2020). We present the results in Table 2.
It is evident that SimpleSBDD–$\mathcal{PO}$ is significantly more efficient than all the other methods. Note that these baselines are not only the best performing, but also the fastest according to (Fu et al., 2022). What makes them much slower than SimpleSBDD are the repeated evaluations of the Vina software, which is approximately 20000x slower than e.g. computing QED (20s vs 0.001s). Furthermore, the only method outperforming SimpleSBDD in terms of Vina score, AutoGrow4, is significantly worse in terms of QED. SimpleSBDD offers an attractive tradeoff between predicted binding affinity and the predicted quality of the generated drug candidates at a much lower computational cost.
Table 2: SimpleSBDD is significantly faster than optimization-based methods. Comparison with optimization-based methods. $\dagger$ denotes method run exclusively on CPU.
@@
SBDD is a problem of generating novel drugs likely to bind to a given target pocket. However, an equally important application is drug repurposing, a problem of investigating new therapeutic uses for existing molecules, which is closely related to determining whether a drug can effectively bind to a specific protein target. This is a significant application since, unlike samples generated by machine learning models, existing drugs have experimentally verified properties and their synthesizability is already established.
@@
To showcase the practicality and efficiency of our scoring model in the context of drug repurposing, we perform the following experiment. For each protein pocket, we randomly sample 16384 molecules from a database and select 100 of them with predicted Vina scores between 5th and 10th percentile. We then compare the average binding affinity and other molecular properties with the reference molecules of the target protein. We found that for each protein pocket it takes around 2 seconds to find 100 diverse molecules with much higher binding affinity and improved other molecular metrics. Furthermore, we repeated the experiment with ChEMBL dataset (Mendez et al., 2018; Davies et al., 2015) and found that $g_{\theta}$ generalizes well across molecular databases. See Appendix J for details.
6 Conclusion, Broader Impact, and Limitations
We advocate a shift in focus from mere model expressivity to robust generalization. Specifically, we bring to the forefront inherent limitations in expressivity of GNNs and prove that they carry over to the multibody domain of SBDD. Furthermore, we show that with a novel model decomposition and explicit binding optimization built into the model, we can outperform state of the art with significantly simplified models. The unprecedented efficiency and success of SimpleSBDD not only highlights the potential to reshape the approach to SBDD but also paves the way for streamlining docking software, especially when assessed with widely adopted tools like Vina and Gnina.
@@ 6 Conclusion, Broader Impact, and Limitations
The very nature of SimpleSBDD as a generative model, especially with its property optimization capabilities, brings with it risks. As SBDD models become increasingly adept at designing molecules with specific properties, we must remain vigilant of the potential unintended outcomes, since streamlining the design process could accelerate the design of harmful biochemicals.
Reproducibility statement
All our experiments are reproducible. Code and the trained models are available under MIT Licence at github.com/rafalkarczewski/SimpleSBDD.
@@
Appendix A Decoupling the unlabelled molecular graph from atom types
@@ A.1 Vina score
Vina uses a scoring function, which approximates the experimentally measured binding affinity with several types of interactions between the protein and ligand atoms [Trott and Olson, 2010]:
@@
Vina uses the BFGS optimiser [Fletcher, 2000]. Due to the redocking procedure (9), the Vina score can be interpreted as "binding potential", i.e. molecule’s binding affinity in its optimal pose.
@@ A.2 Influence of unlabelled molecular graphs on Vina scores
Equation (9) indicates that if Vina employed an ideal optimizer to compute the global minimum, the initial 3D configuration of the molecule would not influence the outcome. Yet, because Vina utilizes an approximate gradient-based optimization method, variations in the initial 3D configuration can indeed affect the Vina score. To measure that, we randomly selected 1,000 protein-ligand complexes from the CrossDocked2020 dataset [Francoeur et al., 2020] and computed the Vina score both before and after modifying the ligand. Here, by ’modifying the ligand’, we refer to replacing its 3D configuration with a randomly generated conformer using RDKit. As expected, a high $R^{2}=0.94$ emerged indicating insignificant impact of the ligand’s initial 3D configuration on the Vina score (See Figure 6(a)).
Furthermore, we experimented with changing ligand’s atom types. For a fixed unlabelled graph $U^{M}$, we re-sample random atom assignments $\mathbf{a}^{M}$ from the empirical atom distributions, while avoiding valence violations (See Appendix A.3). This procedure changes the molecule and therefore its 3D configuration is most likely no longer energetically valid. We therefore use RDKit to generate a random conformer of the new molecule and use it to replace the original 3D coordinates. Surprisingly, the $R^{2}$ coefficient remains relatively high at $R^{2}=0.76$ (see Figure 6(b)). These findings suggest that using just the unlabelled graph of the ligand, one can estimate the Vina score with considerable accuracy. We use this observation in the definition of the model.
A.3 Sampling valid molecules for a given unlabelled graph
Here we provide details on how to randomly modify atom types conditioned on a molecular graph’s structure to obtain valid molecules. Simply assigning random atom types independently to each node usually yields invalid molecules due to violated valence constraints. Whether assigning a specific atom type to a given node $v^{M}_{i}$ results in a violated valence constraint depends on how many and what kind of bonds are connected to it. Specifically, for a bond $e_{ij}$, let $o(e_{ij})$ denote its multiplicity, i.e. $o(e_{ij})=1$ for a single bond, $o(e_{ij})=2$ for a double bond etc. Let now
$\deg(v_{i}^{M})=\sum_{j:e_{ij}\in U^{M}}o(e_{ij})$ (10)
denote $i$-th node’s degree. Node’s degree determines what atom types satisfy valence constraints for that node. It is therefore sufficient to determine the degrees of all nodes in the unlabelled graph and independently sample atom types for each node conditioned on its degree. Formally, we define the distribution over atom types given the molecule’s unlabelled graph as
$p(\mathbf{a}^{M}|U^{M})\coloneqq\prod_{i=1}^{N_{M}}p_{\mathrm{emp}}(a_{i}^{M}|\deg(v_{i}^{M})),$ (11)
where $p_{\mathrm{emp}}(a_{i}^{M}|\deg(v_{i}^{M}))$ is the empirical distribution of atom types for a given degrees of the nodes estimated using the CrossDocked2020 dataset [Francoeur et al., 2020]. This allows for efficient sampling of valid molecules with a fixed unlabelled graph $U^{M}$ so that the distribution of atom type counts matches the empirical one.
@@ A.3 Sampling valid molecules for a given unlabelled graph
A.4 Justification of the model decomposition - Gnina
@@
In section 4.1, we justified our model decomposition choice through empirical evidence indicating that the unlabelled molecular graph inherently contains substantial information, as reflected in the Vina score, even when the ligand’s initial 3D configuration and atom type assignment are varied. We now show that a similar phenomenon can be observed when, instead of Vina, a reportedly significantly more accurate binding software Gnina [McNutt et al., 2021] is used.
We repeat the experiments from the section A.2 with Gnina instead of Vina and find that for changing the initial 3D configuration, there is even smaller impact indicated by $R^{2}=0.97$ (due to Gnina also performing redocking). For changing atom type assignments, $R^{2}=0.76$ as it was for Vina. Please see Figure 7 for details.

### v2

@@ 4.3 Atom sampler
To define the generative model over the atom types given unlabelled structure, we adapt an existing generative model MoFlow (Zang and Wang,, 2020), consisting of an unlabelled graph model $f_{U}$ and a conditional atom model $f_{\mathbf{a}|U}$. The atom model induces a conditional distribution $p_{f_{\mathbf{a}|U}}(\mathbf{a}^{M}|U^{M})$, which we use as our atom sampling distribution:
@@
Table 1: SimpleSBDD finds high-quality drug candidates up to 1000x faster than competing methods. Properties of CrossDocked2020 data and generation methods. $\dagger$ denotes method run exclusively on CPU. "-" denotes that a metric was not reported and we were unable to estimate. Error bars correspond to the standard deviation across test protein pockets. See Appendix H for comparison details.
@@ 5 RESULTS
We follow Peng et al., (2022); Schneuing et al., (2022) and use the CrossDocked2020 dataset (Francoeur et al.,, 2020) for evaluating the models. We use the same train-test split, which separates protein pockets based on their sequence similarity computed with MMseqs2 (Steinegger and Söding,, 2017) and contains 100,000 train and 100 test protein-ligand pairs.
@@ 5.1 Assessing general properties of sampled molecules
We follow the evaluation procedure described by Peng et al., (2022). For each of 100 test protein pockets, we sample 100 molecules. As evaluation metrics we report 1) Vina Score estimating binding affinity (Alhossary et al.,, 2015), 2) High Affinity, which is the percentage of generated molecules with binding affinity at least as good as ground truth, 3) QED, a quantitative estimate of druglikeness, 4) SA, synthetic accesibility, 5) Diversity defined by average pairwise Tanimoto dissimilarity for the generated molecules in the pocket, 6) Novelty, defined by average Tanimoto dissimilarity to the most similar molecule from training set, 7) #Params, the number of trainable parameters and 8) Time to generate 100 molecules in seconds. We include more molecular metrics in Appendix G. As baselines, we use eight recent methods for SBDD: Pocket2Mol (Peng et al.,, 2022), DiffSBDD (Schneuing et al.,, 2022), TargetDiff (Guan et al., 2023a, ), DecompDiff (Guan et al., 2023b, ), FLAG (Zhang et al.,, 2022), DrugGPS (Zhang and Liu,, 2023), D3FG (Lin et al.,, 2024) and EQGAT-diff (Le et al.,, 2024).
@@ 5.2 Additional evaluation
We have demonstrated that SimpleSBDD is very competitive on the standard evaluation criteria for SBDD models. We now discuss its broader applications.
@@
Due to our novel decomposition of the model into the unlabelled graph and atom labels, we can optimize the estimated binding affinity and other properties simultaneously. This is especially important, since molecules without desirable drug-like properties are of diminished practical utility even if the predicted binding is strong. Specifically, given the target protein we first generate an unlabelled graph with strong predicted binding affinity. As we noted in Section 4.2 and elaborated on in Appendix C, there is large diversity in the chemical properties of molecules that share the unlabelled graph. Therefore, for the sampled unlabelled graph, there is enough flexibility in atom types to enable molecular property optimization.
As an illustrative example, we follow Gómez-Bombarelli et al., (2018) and choose $5\cdot\mbox{QED}+\mbox{SA}$ as the property mixture to optimize alongside predicted binding affinity. Concretely, 50 proposals are sampled using the atom type model defined in Equation (7) and the candidate with the best property values is selected. We denote this model variant SimpleSBDD–$\mathcal{PO}$. With this additional procedure we can significantly improve all molecular metrics without sacrificing binding affinity whilst remaining an order of magnitude faster than baselines. See Appendix I for more details.
@@
So far we have compared our method with multiple baselines which generate molecules conditioned on a target protein. Neither SimpleSBDD nor any of these baselines has access to the Vina software during generation. We now also compare with optimization-based methods, i.e. algorithms, which generate molecules by explicitly using Vina during sampling to directly optimize it. Specifically, we compare SimpleSBDD–$\mathcal{PO}$ with the three best performing methods, i.e. Reinforced Genetic Algorithm (Fu et al.,, 2022), 3D-Monte Carlo Tree Search (3D-MCTS) (Du et al.,, 2023) and AutoGrow4.0 (Spiegel and Durrant,, 2020). We present the results in Table 2.
It is evident that SimpleSBDD–$\mathcal{PO}$ is significantly more efficient than all the other methods. Note that these baselines are not only the best performing, but also the fastest according to (Fu et al.,, 2022). What makes them much slower than SimpleSBDD are the repeated evaluations of the Vina software, which is approximately 20000x slower than e.g. computing QED (20s vs 0.001s). Furthermore, the only method outperforming SimpleSBDD in terms of Vina score, AutoGrow4, is significantly worse in terms of QED. SimpleSBDD offers an attractive tradeoff between predicted binding affinity and the predicted quality of the generated drug candidates at a much lower computational cost.
Table 2: SimpleSBDD is significantly faster than optimization-based methods. Comparison with optimization-based methods. $\dagger$ denotes method run exclusively on CPU. Error bars correspond to the standard deviation across test protein pockets.
@@
SBDD is a problem of generating novel drugs likely to bind to a given target pocket. However, an equally important application is drug repurposing, a problem of investigating new uses for existing molecules, which is closely related to determining whether a drug can effectively bind to a specific protein target. This is a significant application since, unlike samples generated by machine learning models, existing drugs have experimentally verified properties and their synthesizability is established.
@@
To showcase the practicality and efficiency of our scoring model in the context of drug repurposing, we perform the following experiment. For each protein pocket, we randomly sample 16384 molecules from a database and select 100 of them with predicted Vina scores between 5th and 10th percentile. We then compare the average binding affinity and other molecular properties with the reference molecules of the target protein. We found that for each protein pocket it takes around 2 seconds to find 100 diverse molecules with much higher binding affinity and improved other molecular metrics. Furthermore, we repeated the experiment with ChEMBL dataset (Mendez et al.,, 2018; Davies et al.,, 2015) and found that $g_{\theta}$ generalizes well across molecular databases. See Appendix J for details.
6 CONCLUSION AND BROADER IMPACT
We advocate a shift in focus from mere model expressivity to robust generalization. Specifically, we bring to the forefront inherent limitations in expressivity of GNNs and prove that they carry over to the multibody domain of SBDD. Furthermore, we show that with a novel model decomposition and explicit binding optimization built into the model, we can outperform state of the art with significantly simplified models. The unprecedented efficiency and success of SimpleSBDD highlights the potential to reshape the approach to SBDD but also paves the way for streamlining docking software, especially when assessed with widely adopted tools like Vina and Gnina.
@@ 6 CONCLUSION AND BROADER IMPACT
The very nature of SimpleSBDD as a generative model, especially with its property optimization capabilities, brings with it risks. As SBDD models improve at designing molecules with specific properties, we must remain vigilant of the potential unintended outcomes, since streamlining the design process could accelerate the design of harmful biochemicals.
7 LIMITATIONS
We acknowledge limitations of current docking-based approaches in real-world drug discovery. While docking scores are a widely used metric in SBDD, they should be interpreted with caution due to the following concerns:
Docking Score Reliability: A high docking score does not necessarily guarantee success in subsequent molecular simulations or experimental validations. Docking scores indicate the predicted binding pose of a ligand to a protein target and should not be overinterpreted as definitive indicators of biological activity.
Dataset Curation and Bias: The quality and curation of datasets can significantly impact docking outcomes. Protein structures from the Protein Data Bank (PDB) often exist as complexes with specific ligands. It is not uncommon to delete the ligand from the complex and dock new molecules instead. This can influence the docking results.
Biological Context: Docking scores do not distinguish between different functional outcomes of binding, such as whether a molecule acts as an inhibitor or an agonist. This limits their utility in predicting the nuanced biological effects of candidate molecules
Correlation with Experimental Data: Docking scores and experimental binding affinities are not always strongly correlated, and docking methods may lack proper calibration against real-world biochemical and biophysical data.
Binding Site Challenges: The binding site of a protein target might be suboptimal or unsuitable for a candidate ligand, further complicating the interpretation of docking results.
Such limitations underscore the need to treat docking scores with caution, and provide motivation for our work – we urge the community to critically evaluate the existing paradigm and redirect the efforts for SBDD (saving the computation expense and time using accelerated approaches such as the one we advocate, and paying additional attention to more reliable investigations such as molecular simulations).
Acknowledgments
This work was supported by the Finnish Center for Artificial Intelligence (FCAI) under Flagship R5 (award 15011052). VG also acknowledges the support from Saab-WASP (grant 411025), Academy of Finland (grant 342077), and the Jane and Aatos Erkko Foundation (grant 7001703). RK thanks Paulina Karczewska for her help with preparing figures.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] Section 4
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] Table 1
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] We include code to reproduce our results as part of the submission.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] Proposition 2 and Lemma 3
    2. (b)
      Complete proofs of all theoretical results. [Yes] Appendix K
    3. (c)
      Clear explanations of any assumptions. [Yes] Section 3
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] The evaluation procedure is described in Section 5 with more details in Appendix H and the submitted code in the supplementary can be used to reproduce our results.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] Data is described in Section 5 and model details are provide in Appendices D and F
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] Tables 1, 2, 4, 5 and 7
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] Appendix D.1 and F
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] The datasets CrossDocked2020, ZINC250k and ChEMBL datasets are properly cited. The MoFlow model that we used is also properly cited.
    2. (b)
      The license information of the assets, if applicable. [Yes] Appendix M
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes] A newly curated dataset described in Section 4 is included in the submitted supplementary material.
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Acknowledgments
Appendix A DECOUPLING THE UNLABELLED MOLECULAR GRAPH FROM THE ATOM TYPES
@@ A.1 Vina score
Vina uses a scoring function, which approximates the experimentally measured binding affinity with several types of interactions between the protein and ligand atoms (Trott and Olson,, 2010):
@@
Vina uses the BFGS optimiser (Fletcher,, 2000). Due to the redocking procedure (9), the Vina score can be interpreted as "binding potential", i.e. molecule’s binding affinity in its optimal pose.
@@ A.2 Influence of unlabelled molecular graphs on Vina scores
Equation (9) indicates that if Vina employed an ideal optimizer to compute the global minimum, the initial 3D configuration of the molecule would not influence the outcome. Yet, because Vina utilizes an approximate gradient-based optimization method, variations in the initial 3D configuration can indeed affect the Vina score. To measure that, we randomly selected 1,000 protein-ligand complexes from the CrossDocked2020 dataset (Francoeur et al.,, 2020) and computed the Vina score both before and after modifying the ligand. Here, by “modifying the ligand", we refer to replacing its 3D configuration with a randomly generated conformer using RDKit. As expected, a high correlation at $\rho=0.97$ emerged indicating an insignificant impact of the ligand’s initial 3D configuration on the Vina score (See Figure 5(a)).
Furthermore, we experimented with changing ligand’s atom types. For a fixed unlabelled graph $U^{M}$, we re-sample random atom assignments $\mathbf{a}^{M}$ using MoFlow (Equation 7). This procedure changes the molecule and therefore its 3D configuration is most likely no longer energetically valid. We therefore use RDKit to generate a random conformer of the new molecule and use it to replace the original 3D coordinates. Surprisingly, the correlation remains relatively high at $\rho=0.83$ (see Figure 5(b)). These findings suggest that using just the unlabelled graph of the ligand, one can estimate the Vina score with considerable accuracy. We use this observation in the definition of the model.
@@
A.3 Justification of the model decomposition - Gnina
@@
In section 4.1, we justified our model decomposition choice through empirical evidence indicating that the unlabelled molecular graph inherently contains substantial information, as reflected in the Vina score, even when the ligand’s initial 3D configuration and atom type assignment are varied. We now show that a similar phenomenon can be observed when, instead of Vina, a reportedly significantly more accurate binding software Gnina (McNutt et al.,, 2021) is used.
We repeat the experiments from the section A.2 with Gnina instead of Vina and find that for changing the initial 3D configuration, there is even smaller impact indicated by $\rho=0.99$ (due to Gnina also performing redocking). For changing atom type assignments, $\rho=0.85$ slightly higher than it was for Vina. Please see Figure 6 for details.
A.4 Generality of the findings
We have demonstrated that significant portion of binding information, as measured by binding software, is contained in the unlabelled molecular graph. So far, we performed all experiments on the gold-standard dataset for SBBD - CrossDocked2020. We now repeat the analysis from section 4.1 for PDBind (Liu et al.,, 2015), specifically its refined subset containing $\sim 5000$ highest-quality protein-ligand complexes. Arguably, this is a better benchmark than CrossDocked2020, which has been augmented with protein-ligand complexes, without experimentally measured binding (the "cross-docking" procedure).
We performed the same experiment as in Section 4.1, i.e. measuring the impact of 1) changing the initial 3D configuration, and 2) changing atom types. In both cases, we measured the binding with both Vina and Gnina. We observed that the correlation is slightly lower, but still large enough to infer that a significant portion of the binding information is contained in the unlabelled graph. This suggests that our findings hold more broadly across different datasets, including those compiled from experimental measurements. Please see Table 3 for more details.
Table 3: Unlabelled molecular graph contains a significant portion of binding information. Findings hold across different binding software and datasets.
Experiment Software CrossDocked2020 PDBind
Changing atom Vina 0.827 0.781
types Gnina 0.849 0.797
Changing intitial 3D Vina 0.971 0.953
configuration Gnina 0.989 0.979